Wednesday, February 07, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

Blue Jeans Connection Information

  1. Enter Meeting ID: 654704307
  2. Press #


  1. Review systems owners and update information if needed
  2. Point system owners to Wiki's that need to be addressed for the upcoming commissioning run
  3. Address system specific points of interest for the upcoming commissioning run

System Owners





General Questions to be Addressed by Individual Groups

  1. What is the current status on the Hardware?
    • Identify any remaining bad detectors/channels. (Show a quick slide that indicates status?)
    • Are there plans to repair them prior to the KPP run, or after KPP?
  2. What is the status of the Software associated with your detector?
    • Have you successfully run the analyzer and been able to identify signals for your detector?
    • Have the HV <-> ADC <-> TDC channel mappings all been verified to be correct?
    • Have any relevant survey/position information been updated your detector's PARAM files?
    • Have all calibration constants been updated in the PARAM files (ie. gains, time offsets, etc...)?
    • If beam is needed to finalize these params, is the software chain ready to go to generate them quickly?
  3. Can you think of anything else that needs to be completed before we take beam?
    • Missing software you need?
    • Missing survey/position information to complete the PARAM files?

Specific Questions to be Addressed

  1. HMS & SHMS Drift Chambers
    • Can we acquire tracks from the cosmic data yet?
  2. NGC
    • Are the HV signal cables hooked up?



  • All groups are requested to update documentation on the wiki link
    • We will use these as references for shift crew later.
    • See links under System Owners
    • Use the HMS Drift Chambers entries as an example
    • Links should point to documents in the DocDB
    • Don't just upload to the wiki.
  • Set HW limit on all HV cards to set a reasonable hardware limit
    • Dave Mack volunteered
      • Brad will get him a screwdriver


Drift Chambers

  • System Owner: E. Pooser
  • Ethane:Argon now in use
  • Working on final offsets for alignment params from Survey added to PARAM files
    • Tracking is working though
  • A few dead channels(known), no hot channels
  • Calibration constants are up-to-date in PARAM files
  • Need to recheck deficiencies with the Eth:Ar gas (should happen this week)


  • System owners: S. Malace, M. Jones]
  • Hardware in good shape
  • Rough timing calibration is done
  • Generally in good shape


  • System Owners: H. Mkrtchyan, Yerevan
  • PMTs are powered up, but can't see good signals now
    • Saw signals in the past: see Simona's ERR talk
    • Needs to be followed up -- not sure what changed
      • Maybe need to check the gains / HVs

Heavy Gas Cherenkov

  • System owners: B. Sawatzky, E. Pooser
  • Pump and fill with CO2 is in progress now
    • Hopefully complete by end of week
  • Signals cables have been patched through from the Hut to the F250's upstairs
  • PMTs are showing dark current signals
    • They need to be reconnected to the patch panel in the hut


Drift Chambers

  • System owners: E. Christy, D. Biswas
  • Need to get resurveyed (Walter)
  • No tracks from the analyzer yet
    • Jure is trying to figure out the PARAM file settings
  • Ethane:Argon now in use
  • 3 cards of missing data
  • Plan to do voltage scan and check efficiency with new Eth:Ar mix
  • Need to do drift times and T0 calibration in the software


  • System Owners: S. Malace, M. Jones]
  • Plastic scintillators in good shape
  • Gain matching complete
  • Quartz plane status
    • Bars 7--16 (middle) are the 'working' bars for now (though bar 16 is poor)
    • 4 of the 9 bars show correlated signal using using course tracking cuts from hodoscopes
    • Real tracking from SHMS chambers will help this analysis when available
    • Another 4 bars are questionable, playing with gains, etc
    • 9th bar (#16) has XP tube on one end which in really bad shape


  • System owners: H. Mkrtchyan, Yerevan
  • Preshower
    • All channels good
    • Gains are high, can see muons in all channels
    • May need to remove amplifier and replace with passive splitter
      • Be sure to communicate with Brad before doing this
  • Shower
    • One bad PMT (pulling too much current)
    • Block is outside acceptance, so maybe fix in summer?
    • Very small signals visible in most channels, could be due to backwards cosmics?
      • Not sure if this is OK or not

Aerogel Cerenkov

  • System owners: A. Mkrtchyan, T. Horn
  • All signals present in DAQ
  • Large attenuation noted between upstairs and downstairs
  • Seems to be confusion about summing, splitting upstairs, need to be clarify what we really want -- talk to Brad
  • One missing channel, Arthur working on it
  • Most channels seem to be OK downstairs but show a lot of noise Counting House, this needs to be solved.


  • System owners: D. Day, E. Pooser
  • Tank will be filled with Argon after it is installed
  • Waiting for installation into the SHMS stack (presently on the Hall C floor)
  • HV and signal cables are run and verified into DAQ upstairs
  • Need to remove paper PMT targets
    • Donal will organize for this to happen next week.


  • System owners: G. Huber, B. Li
  • Will be filled with CO2
    • Pump and fill is in progress this week (Brad)
  • One dead channel chased to bad cable -- fixed
  • All channels look alive
  • Hardware and basic software set up is in good shape
  • May want to switch FADCs front-end digitizer FSR to 2V later